Aqueous hydrobromic acid (HBr) will react with solid sodium hydroxide (NaH) to produce aqueous sodium bromide (NaBr) and liquid water (H, 0).

Suppose 75.2 g of hydrobromic acid is mixed with 14. g of sodium hydroxide. Calculate the maximum mass of sodium bromide that could be produced by the

chemical reaction. Be sure your answer has the correct number of significant digits.


1
Expert's answer
2022-04-05T09:10:06-0400

HBr + NaOH = NaBr + H2O

n(HBr) = m(HBr) / Mr(HBr) = 75.2 g / 80.91 g/mol = 0.93 mol

n(NaOH) = m(NaOH) / Mr(NaOH) = 14.0 g / 40.0 g/mol = 0.35 mol

NaOH is limiting reactant

n(NaBr) = n(NaOH) = 0.35 mol

m(NaBr) = n(NaBr) × Mr(NaBr) = 0.35 mol × 102.894 g/mol = 36.0 g